### v2

# The Changes of the Northern Hadley Cell Strength in Reanalyses and Radiosonde Observations

Abstract
This study examines mean meridional winds and their trends in the Northern Hadley Cell (NHC) from 1980 to 2022 using reanalysis datasets and radiosonde observations. Compared to radiosonde data, reanalyses underestimate the mean upper-tropospheric poleward flow of the NHC but accurately capture the mean equatorward flow in the lower troposphere. While climate models generally project a weakening of the NHC, our study finds no significant trend in radiosonde observations, adding to the uncertainty in future climate projections. In contrast, reanalyses indicate a strengthening, primarily due to an intensification of the upper-tropospheric poleward flow. Our examination of ERA5 analysis increments confirms that the NHC strengthening trend in ERA5 is not an artefact of data assimilation. Instead, the increments correct the first-guess, which underestimates the strength of the NHC, nudging it toward a stronger circulation.
Plain Language Summary
The Hadley circulation (HC) is a key atmospheric system that transports heat and moisture between the tropics and subtropics. It consists of two cells that converge near the equator, characterized by rising air in the tropics, poleward flow in the upper troposphere, descending air in the subtropics, and lower-tropospheric trade winds returning toward the equator. This study investigates changes in meridional winds within the Northern Hadley cell (NHC) and its overall strength using weather balloon wind measurements (radiosondes) and reanalysis datasets, which reconstruct past weather conditions. While climate models project a weakening of the NHC, supported by studies based on sea-level pressure data, reanalyses suggest that the NHC has strengthened over recent decades. However, our study finds a neutral trend in NHC strength based on radiosonde data, which further contributes to the uncertainty and potential unreliability of future climate projections. Although previous research suggested that NHC strengthening in reanalyses results from data assimilation artefacts, this study shows that in ERA5, data assimilation corrects the weak NHC bias, but does not drive its strengthening trend.
1 Introduction
The Hadley circulation (HC) is a large-scale thermally driven atmospheric meridional overturning circulation that spans tropical and subtropical regions. It consists of two cells extending from the Inter-tropical convergence zone (ITCZ) to the subtropics and plays a pivotal role in the transport of moisture, heat and momentum between the tropics and higher latitudes (e.g. Peixoto and Oort, 1992).
Recent studies have highlighted conflicting historical trends in Northern Hadley cell (NHC) strength between climate models and reanalyses (Chemke and Polvani, 2019). While reanalyses generally suggest a strengthening of the NHC in recent decades (Tanaka et al., 2004; Mitas and Clement, 2005; Sohn and Park, 2010; Stachnik and Schumacher, 2011; Nguyen et al., 2013; Pikovnik et al., 2022; Zaplotnik et al., 2022), climate model simulations depict a weakening of the NHC over the same period (Vallis et al., 2015; Hu et al., 2018; Xia et al., 2020). This disparity introduces uncertainty into our understanding of the atmospheric circulation response to climate change and, consequently, the future impact of the Hadley circulation on tropical and subtropical precipitation. The state-of-the-art ERA5 reanalysis (Hersbach et al., 2020) also indicates a strengthening of both Hadley cells in recent decades (Zaplotnik et al., 2022; Pikovnik et al., 2022). However, the strengthening in the NHC is observed only until around the year 2002, followed by a mild decline after that.
This inconsistency necessitates verification against direct atmospheric observations to better understand the true historical changes in NHC strength. For example, Chemke and Yuval (2023) recently demonstrated that in climate models the meridional gradient of zonal-mean sea-level pressure between the tropics and subtropics (denoted SLPy) can be used as a proxy for Hadley cell strength. They compared the time series of SLPy, derived from reanalysis data, historical climate model runs, and observation-based ICOADS R3.0 sea-level pressure (SLP) data (Freeman et al., 2017). Their findings showed that climate models closely align with the surface proxy for the HC strength, suggesting a significant weakening of the NHC in recent decades. In contrast, most reanalyses indicated an NHC strengthening or neutral trend, except those that only assimilate conventional surface observations (e.g., NOAA 20CRv3, ERA 20C). This discrepancy between observations and reanalyses that assimilate satellite observations was speculated to arise from artefacts in the assimilation of satellite observations in reanalyses (Chemke and Polvani, 2019).
Although surface-based proxies for Hadley circulation strength are invaluable, describing circulation using long-term records of reliable and accurate upper-air meridional wind data is crucial. These measurements capture the entire circulation directly, rather than describing only a portion (e.g., the lower branch) of the Hadley cell. This study compares NHC strength in modern reanalyses and radiosonde observations and examines the impact of data assimilation systems in the ERA5 reanalysis on changes in NHC strength.
The study is organized as follows. Section 2 describes the data and methodology, Section 3 presents the results, while a discussion, conclusions, and outlook are given in Section 4.
2 Methodology
2.1 Data
This study utilizes reanalyses data from ERA5 (Hersbach et al., 2020), JRA-55 (Kobayashi et al., 2015), MERRA-2 (Gelaro et al., 2017), and NOAA 20CRv3 (Slivinski et al., 2021) to study the strength of the Northern Hadley cell. ERA5 features the highest model resolution, the most detailed model physics, and the largest volume of assimilated data (Hersbach et al., 2020). This results in a good agreement with observations of tropospheric temperature, wind, humidity, and precipitation (Simmons, 2022). ERA5 also notably reduces the biases in the surface meridional wind and the horizontal wind divergence over the oceans (Rivas and Stoffelen, 2019). Additionally, the quality of the atmospheric energy, moisture and mass budgets derived from ERA5 is demonstrably better than the previous generation reanalyses, such as ERA-Interim (Mayer et al., 2021). The Japanese 55-year Reanalysis (JRA-55) has a larger warm bias in the upper troposphere, a notable imbalance in global mean net energy fluxes at the top and bottom of the atmosphere, and excessive precipitation over the tropics (Kobayashi et al., 2015; Harada et al., 2016). The Modern-Era Retrospective Analysis for Research and Applications, Version 2 (MERRA-2) has a significant positive temperature bias in the mid-to-upper troposphere and excessive precipitation in regions of high tropical topography (Gelaro et al., 2017). More details of the respective reanalyses can be found in Fujiwara et al. (2022).
We used NOAA 20CRv3 to evaluate the impact of satellite observations on HC strength since the NOAA reanalysis does not assimilate satellite data but relies only on sea-level pressure data, along with sea surface temperature boundary conditions. Its long time series and multidecadal averages of mass, circulation, and precipitation fields closely align with other modern reanalyses, as well as with station- and satellite-based products. However, substantial biases in temperature and wind above 300 hPa suggest a potential inadequacy for upper atmosphere studies in reanalysis that are based solely on surface pressure assimilation (Slivinski et al., 2021).
The meridional wind ($v$) is provided at standard pressure levels: 37 pressure levels in ERA5 and JRA-55, 42 levels in MERRA-2 and 28 levels in NOAA 20CRv3; and on a latitude-longitude grid with 1° resolution in ERA5 and NOAA 20CRv3, 1.25° in JRA-55 and 0.5° x 0.625° in MERRA-2. In the reanalyses, the analysis times are set at 00 and 12 UTC, and their monthly means are compared to the corresponding monthly means of radiosonde observations. The exception is MERRA-2, whose monthly means of daily means are compared to the monthly mean of radiosonde observations at 00 and 12 UTC.
We then compare all reanalyses with an observational dataset, using quality-controlled radiosonde data from the Integrated Global Radiosonde Archive Version 2.2 (IGRA 2.2, Durre et al. (2018)). This study utilizes monthly-mean values of the meridional wind at 21 standard pressure levels ranging from 1000 to 1 hPa. Monthly means are computed for pressure levels at nominal times of 00 UTC and 12 UTC, considering data collected within two hours of each nominal time. The IGRA 2.2 dataset encompasses specific years/months in which a satisfactory number of standard pressure-level observations are obtained at fixed observing stations on land. The vertical distribution of radiosonde observations is relatively uniform throughout the troposphere, except below the 850 hPa level, which lies beneath the surface in many regions (Figure S1). The observation locations are unevenly distributed across the globe (Figure 1a,b), primarily concentrated over continental land areas, with their density generally corresponding to the distribution of the population. Thus, NHC, the focus of this study, is better observed than the Southern Hadley cell (Figure S2).
We analyze 43 years of radiosonde data from 1980 to 2022, and up to 43 years of reanalysis data, covering the period 1980-2022 for ERA5, JRA-55 and MERRA-2, and 1980-2015 for NOAA 20CRv3. Throughout this study, we denote the reanalysis meridional wind as $v_{a}$ and the observation-based meridional wind as $v_{o}$.
Figure 1: Annual-mean meridional wind for (a) 200-300 hPa and (b) 700-850 hPa pressure levels, based on ERA5 data for 1979-2022. In (a,b), the shaded region between the pink lines at 36°S and 38°N represents the approximate maximum extent of the two Hadley cells. Pink dots indicate the locations of radiosonde observations for each pressure level. Histograms along the top and left edges show the latitudinal and longitudinal distribution of the data. (c) displays spherical Voronoi tessellation for radiosonde observations at 850 hPa in May 2002, at 00 UTC. The shaded region between the pink lines at 6°N and 31°N represents the average long-term extent of the annual-mean NHC. The dots mark the locations of all radiosonde observations, with pink dots specifically indicating those whose Voronoi faces (grey polygons) contribute to the mean meridional wind within the NHC boundaries. Their contribution is weighted according to the area of the polygon within those boundaries.
2.2 Computing Analysis Increments from Reanalysis’ Data Assimilation Procedure
Analysis increments are corrections applied to the short-range forecast to produce an optimal estimate of the weather state at a specific time, referred to as the analysis. These corrections are derived through the assimilation of Earth-system observations, a process known as data assimilation. The corrections are applied cyclically in each assimilation window (e.g. every 6 or 12 hours). We evaluated the analysis increments ($\delta\mathbf{x}_{a}$) as the difference between the analysis values (represented by the state vector $\mathbf{x}_{a}$) valid at 00/12 UTC and the previous short-range (6-hour) forecast (first-guess, $\mathbf{x}_{f}$) valid at the same time and initialized at 18/06 UTC, i.e. $\delta\mathbf{x}_{a}=\mathbf{x}_{a}-\mathbf{x}_{f}$. Only the ERA5 dataset includes publicly accessible first-guess (forecast) data, enabling the calculation of analysis increments and first-guess departures.
2.3 Comparing Reanalyses with Observations
Reanalyses and irregular observations can be directly compared in the observation space by computing the analysis departures ($\mathbf{d}_{a}$). These are determined by contrasting radiosonde observations, represented by the observation vector $\mathbf{y}$, with analysis values mapped to the observation space using the observation operator $\mathcal{H}$, i.e. $\mathbf{d}_{a}=\mathbf{y}-\mathcal{H}(\mathbf{x}_{a})$. The observation operator $\mathcal{H}$ generates model equivalents of the observations by linearly interpolating the analysis values to the observation points. Analogously, first-guess departures ($\mathbf{d}_{f}$) are computed as the differences between observations and the short-range forecasts from the reanalysis model, i.e. $\mathbf{d}_{f}=\mathbf{y}-\mathcal{H}(\mathbf{x}_{f})$. In this study, $\mathbf{x}$ represents a set of meridional wind fields at various pressure levels combined into a single state vector, while $\mathbf{y}$ denotes a set of monthly-mean observations of the meridional wind. Because observation data are provided at 00 and 12 UTC, the corresponding monthly means of forecasts (first-guesses), analyses, and analysis increments are computed separately for 00 UTC and 12 UTC.
2.4 Hadley Circulation Strength
The Hadley circulation strength is typically evaluated using the mass-weighted zonal-mean stream function $\psi$ in the latitude-pressure ($\varphi$-$p$) plane (Pikovnik et al., 2022). A stream function $\psi$ can be computed using the zonally-averaged mass conservation equation (Oort and Yienger, 1996), typically by the vertical integration of the zonal-mean meridional wind $[v]$ as
$\psi(\varphi,p)=(2\pi R\cos\varphi/g)\,\int_{0}^{p}[v](\varphi,p^{\prime})\,\mathrm{d}p^{\prime}\,,$ (1)
where $R$ is Earth’s radius, $g$ is gravity, $\varphi$ is latitude, and $p$ is pressure. The widely used HC strength metrics are based on single-point values (maxima or minima) of $\psi$ at a certain level or within the whole Hadley cell (Pikovnik et al., 2022). Accurately computing the zonal-mean meridional wind requires data that are regularly spaced in the periodic zonal domain, as found in reanalysis or climate model outputs. To obtain the zonal average from non-uniform observational data, we compute a weighted average of the meridional wind values $v_{i}$ within each latitude belt, as follows
$[v]_{V}(\varphi,p)=\left(\sum_{i}w_{i}v_{i}\right)/\left(\sum_{i}w_{i}\right)\,.$ (2)
The weight $w_{i}$ assigned to each data point corresponds to the area of its associated Voronoi cell within a given latitude belt. Figure 1c illustrates how Voronoi cells are used to compute the mean meridional wind over the entire spherical region bounded by the long-term annual-mean NHC extent, denoted as $[v]_{V}^{\mathrm{NHC}}$. The Voronoi cells are generated using spherical Voronoi tessellation (Caroli et al., 2009) based on radiosonde data at each pressure level and time instance. Polygon edges intersecting NHC boundaries are adjusted prior to computing the spherical area of each cell. We denote Voronoi-based averaging as $[\cdot]_{V}$. For a fair comparison, the same method is applied to reanalysis data by first interpolating the reanalysis values to observation locations ($\mathcal{H}(v_{a})$), followed by Voronoi averaging.
By meridionally averaging the zonal-mean stream function across the Northern Hadley cell, the stream function can be approximated as
$\tilde{\psi}(p)=\langle\psi(\varphi,p)\rangle\approx(2\pi R\,\langle\cos{\varphi}\rangle/g)\,\int_{0}^{p}[v]_{V}^{\mathrm{NHC}}(p^{\prime})\,\mathrm{d}p^{\prime}\,,$ (3)
[…]
$\overline{\tilde{\psi}}=\left(\sum_{k=1}^{N}\tilde{\psi}_{k}\Delta p_{k}\right)/\left(\sum_{k=1}^{N}\Delta p_{k}\right)\,,$ (4)
where $N$ is the number of used pressure levels $p_{k},k=1,\ldots,N$, level $p_{0}$ corresponds 70 hPa, level $p_{N}$ to 1000 hPa, and the weights are $\Delta p_{k}=(p_{k+1}-p_{k-1})/2$, and $\psi_{k}=\psi(p_{k})$.
3 Results
In this section, we (1) compare the meridional winds in ERA5 and radiosonde observations, (2) verify the data assimilation in ERA5 against radiosonde observations, and (3) analyze the NHC strength and its trends in observations and different reanalyses.
3.1 Comparison of meridional winds in ERA5 and radiosonde observations
The representation of the Hadley circulation in ERA5 in terms of the meridional wind cross section is shown in Figure 2a, and the stream function representation is shown in Figure S4. On average, the annual-mean Hadley circulation in ERA5 extends from 32°S to 31°N, with the boundary between the Northern Hadley cell (NHC) and the Southern Hadley cell (SHC) located at 6°N (Pikovnik et al., 2022). Figure 2a captures the main features of the Hadley circulation, the upper-tropospheric poleward flow and the lower-tropospheric equatorward flow. Unlike the zonal-mean picture, the ERA5 meridional winds interpolated to the observation space and averaged using the Voronoi tessellation method reveal a somewhat different pattern (Figure 2b). Specifically, the upper-tropospheric poleward flow in NHC extends further north in the zonal mean of the observation space. This discrepancy arises mainly because of the limited coverage of the upper-tropospheric equatorward flow over the Atlantic (see Figure 1a). Despite the concentration of radiosonde sites over the continents, the Voronoi averaging ensures that the zonal-mean in the observation space adequately estimates the zonal-mean meridional winds and does not relate to zonal-mean meridional winds over land (Figure S5).
Figure 2: (a) Annual-mean zonal-mean meridional wind in ERA5 model space, (b) in ERA5 observation space, and (c) in radiosonde data. The zonal averaging of meridional winds in (b,c) is performed using the Voronoi tessellation method. (d) Meridional winds from (a,b,c), meridionally averaged over the extent of the Northern Hadley cell (from 6°N to 31°N) during the 1980-2022 period.
Radiosonde data suggest that the NHC extends even further north (Figure 2c) and that the upper-tropospheric poleward flow is up to 0.5 m/s stronger than its ERA5 observation-space counterpart (Figure 2b). Conversely, the radiosonde data show a somewhat weaker equatorward flow near the surface. In contrast to the NHC, the SHC is poorly sampled by radiosondes. This leads to large discrepancies between the zonal-mean values in the ERA5 model-space and observation-space zonal means, especially in the upper troposphere, where the magnitude of the mean meridional wind is severely underestimated (see Figures 2a and 2b). SHC is therefore not part of our study.
To mitigate the effects of inhomogeneous radiosonde data, a Voronoi averaging of the meridional winds was applied over the annual-mean NHC extent, ranging from 6°N to 31°N. Figure 2d compares the mean meridional wind in the physical space of ERA5 with the mean meridional wind of ERA5 in the observation space of the radiosondes. The results suggest that, despite the limited number of observation points, an advanced weighting method such as Voronoi tessellation can effectively represent the mean meridional wind profile in the physical space much better than other averaging methods we tested (i.e., simple averaging in observation space and binning averaging; Figure S6). This is particularly true for the equatorward flow in the lower troposphere, while the poleward flow in the upper troposphere tends to be overestimated. Nevertheless, the observations clearly show that ERA5 underestimates the poleward flow in the upper troposphere. However, it should be noted that despite Voronoi averaging, perfect closure of the mass fluxes in the observation space was not achieved, particularly for the radiosonde data (see Figures S7 and S8).
3.2 Verification of data assimilation in ERA5 against radiosonde observations
The underestimation of the NHC upper-tropospheric poleward flow in the ERA5 first-guess against observations (Figure S9a) suggests that the data assimilation increments should correct the bias by strengthening the circulation. Indeed, Figure 3a shows that the data assimilation system generates a systematic quadrupole pattern of zonal-mean meridional wind analysis increments. As demonstrated by the stream function of the analysis increments, the data assimilation intensifies both the northern and southern Hadley cell (Figure 3b), and leads to enhanced upper-tropospheric poleward flow and lower-tropospheric equatorward flow in the ERA5 analysis compared to its first-guess (Figure S9a), in agreement with Mayer et al. (2025, their Figure S4).
Figure 3: (a) Annual-mean zonal-mean analysis increments of meridional wind computed in physical space; (b) annual-mean stream function (contours) and the stream function of the analysis increment (colors); (c) analysis increments of meridional wind in the radiosonde observation space; (d) first-guess departures of meridional wind, representing the difference between observations and the short-range forecast before data assimilation. All panels correspond to the ERA5 reanalysis and are averaged over the 1980-2022 period. The zonal averaging in (c,d) is performed using the Voronoi tessellation method. In (b), contour lines represent the annual-mean zonal-mean Hadley circulation strength, with red contours indicating positive stream function values ($0.1$, $0.3$, $0.6$, $1$$\times$ 1011 kg s-1) and blue contours showing their negative counterparts ($-0.1$, $-0.3$, $-0.6$, $-1$$\times$ 1011 kg s-1). Mean increments are represented by color shading.
To better understand this process, we also examined the analysis increments in the radiosonde observation space (Figure 3c) and compared them to the first-guess departures in the same space (Figure 3d). The results show that the data assimilation system corrects the underestimation of the HC strength present in the ERA5 first-guess, which originates from the underlying forecast model, but only partly compensates for it. This provides further evidence that NHC is underestimated in ERA5 reanalysis.
We further investigate whether the data assimilation system in ERA5 reanalysis contributes to the strengthening trend in NHC, as speculated by Chemke and Polvani (2019). First, we observe that the annual-mean zonal-mean first-guess departures of meridional winds have decreased over recent decades, reducing the magnitude of the analysis increments applied by data assimilation (Figures S10, S11c). Although the analysis increments tend to strengthen the circulation in recent decades, their magnitude has been decreasing (Figure S11c). This suggests that the analysis increments counteract the strengthening trend seen in the first-guess (Figure S11b), leading to weaker strengthening of the (re)analysis (Figure S11a).
Since the same forecast model is used throughout the reanalysis period, the reduction in the magnitude of analysis increments likely reflects improvements in the global observing system, leading to progressively more accurate analyses. In turn, this enhances the quality of the first-guess produced by the ERA5 forecast model, reducing first-guess departures and completing the feedback loop. Our research thus suggests that NHC strengthening in the ERA5 reanalysis is not a direct artefact of data assimilation. While our investigation of analysis increments has been restricted to meridional winds, note that data assimilation may also introduce artefacts in other variables, which could drive spurious trends in HC strength in the first-guess.
In addition, the analysis increments slightly shift the annual-mean boundary between the two Hadley cells southwards (Figure 3b). However, this is beyond the scope of this study.
3.3 NHC strength and its trends in reanalyses and observations
Finally, we compared the mean meridional wind in the observation space, Voronoi-averaged over the extent of the NHC, across multiple reanalyses (ERA5, JRA-55, MERRA-2, and NOAA 20CRv3) and observations. Figures 4a,c show that all reanalyses underestimate the upper-tropospheric poleward flow, particularly NOAA 20CRv3. This bias is likely due to the relatively coarse resolution of the reanalysis models, which tend to underestimate convective activity, leading to a weaker divergent outflow in the upper troposphere. Preliminary results suggest that this issue is mitigated when higher resolution forecast models are used in data assimilation cycling. In contrast, the trade winds in the lower troposphere closely match the observations.
Analyzing trends in the mean meridional wind profile, we find that the reanalyses show a significant strengthening of the poleward flow (Figures 4b,d). The trends in the mid-troposphere (500 hPa) and lower levels in the reanalyses also largely align with observations, indicating a weakening of the trade winds. This is consistent with the reported decrease in the zonal-mean meridional gradient of mean-sea-level pressure (Chemke and Yuval, 2023).
To further investigate these changes, we computed the strength of the Northern Hadley cell using stream-function metric (Equations 3 and 4), providing a broader perspective on circulation changes and trends. Figure 4e presents the standardized time series of the NHC strength, which shows a matching variability of the NHC strength in reanalyses and observations since around the 1990s, while there is a mismatch in the 1980s. In the observation space, the reanalyses consistently depict a significant strengthening (Figure 4e), whereas observations show a more nuanced picture, with an insignificant strengthening over the 1980-2015 period and a neutral trend over the 1980-2022 period. However, as shown in Figure 4e, much of the apparent strengthening trend can be attributed to the low NHC strength in the 1980s, which reanalyses also underestimate relative to observations. Similarly to the results reported in Pikovnik et al. (2022) and Zaplotnik et al. (2022), the NHC strengthening in ERA5 continued until around 2002, followed by a mild decline.
Reanalyses, which only assimilate conventional surface observations, such as NOAA 20CRv3, are often speculated to better capture climate trends in tropical circulations. However, our results indicate that while NOAA 20CRv3 accurately represents trends in surface meridional winds, it misrepresents both mean meridional winds and their trends throughout the troposphere. This occurs despite the dataset not being affected by potential artefacts from satellite data assimilation. Given that surface observations alone fail to accurately constrain mid- and upper-tropospheric flow, we question whether surface observation proxies, such as the meridional gradient of the mean sea level pressure, can serve as reliable indicators of Hadley circulation strength and its long-term changes, despite their strong correlation with Hadley circulation strength in climate models (Chemke and Yuval, 2023).
Figure 4: (a,c) Mean meridional wind in the observation space averaged across the Northern Hadley cell (from 6°N to 31°N) using the Voronoi tessellation method. (b,d) Multidecadal trends of the mean meridional wind in (b) ERA5, JRA-55 and MERRA-2 reanalysis, and (d) NOAA 20CRv3 reanalysis compared to observations. (e) Standardized time series of the stream function $\overline{\tilde{\psi}}$ from 1980 to 2022. (f) Multidecadal trend of $\overline{\tilde{\psi}}$. (a,b) are shown for the period 1980-2022, and (c,d) are shown for the period 1980-2015. Shaded areas in (b,d) indicate the standard error of the trend, which depends on the time series length, the correlation coefficient between both time series, and the standard deviations of both time series. The black bars in (f) denote the standard errors of the trend estimate.
4 Conclusions and Outlook
This study evaluates the strength and long-term trends of the Northern Hadley cell (NHC) by comparing meridional wind data from modern reanalyses with radiosonde observations. We first demonstrate that the Voronoi averaging of ERA5 data, interpolated to radiosonde observation locations, captures the key features of the Northern Hadley cell, despite the inhomogeneous and sparse radiosonde network, which is limited to land areas and isolated islands. On the other hand, the radiosonde network in the Southern Hemisphere is too sparse for a meaningful representation of the Southern Hadley cell.
Focusing on the NHC, our assessment of mean meridional winds reveals that the NHC is generally stronger in observations than in reanalyses, primarily due to stronger upper-tropospheric poleward flow. Additional investigation of ERA5 first-guess departures and analysis increments indicates that the ERA5 forecast model systematically underestimates the NHC strength. Data assimilation corrects it by imposing meridional wind increments that strengthen the circulation, confirming that ERA5 has a negative bias in the NHC strength. Our results are consistent with Mayer et al. (2025), who showed that the data assimilation in ERA5 enhances the surface meridional winds over the Pacific. Using a completely different approach, Vonich and Hakim (2025) showed that optimal initial-condition perturbations for improving weather predictability involve large-scale corrections to ERA5, indicative of a stronger Hadley circulation, further suggesting that ERA5 underestimates the strength of the Hadley circulation.
Consistent with previous studies of (Nguyen et al., 2013; Chemke and Polvani, 2019; Zaplotnik et al., 2022; Pikovnik et al., 2022), reanalyses suggest a significant strengthening of the NHC in the recent decades, driven mainly by an intensification of upper-tropospheric poleward flow. Conversely, lower-tropospheric equatorward flow is weakening in the reanalyses, in line with the reported decline in the meridional gradient of mean sea level pressure (Chemke and Yuval, 2023). Contrary to the findings in this study and contrary to climate model projections, our research based on radiosonde observations indicates a neutral trend in NHC strength over the 1980-2022 period and an insignificant strengthening over 1980-2015. Furthermore, we show that the analysis increments do not contribute to the strengthened NHC trend in ERA5.
Our findings (1) raise critical questions about the reliability of reanalyses to estimate long-term HC strength trends, particularly in the context of climate change; (2) suggest that the reanalyses based on surface observations are insufficient to constrain the full vertical extent of atmospheric circulation properly; (3) underscore the essential role of conventional upper-air observations in verifying global circulation changes, particularly over remote ocean areas; (4) raise concerns about the reliability of climate model projections, which indicate a weakening of the NHC.
Overall, this study emphasizes the importance of robust observational datasets in validating reanalysis outputs and refining our understanding of climate dynamics, particularly regarding the future evolution of the Hadley circulation in a changing climate. Strengthening observational networks will be critical for reducing uncertainties in reanalyses and better constraining the long-term behavior of the Hadley circulation. This, in turn, will enhance the verification of climate model simulations and improve confidence in future climate projections.
Acknowledgements
Matic Pikovnik acknowledges funding from the Slovenian Research and Innovation Agency (ARIS) Programme P1-0188 and Grant MR-56969. Žiga Zaplotnik acknowledges the funding by the European Union under the Destination Earth initiative. MP acknowledges funding from the Slovenian National Recovery and Resilience Fund project ULTRA within subproject 6.03 Environmental Technologies for Climate Change Mitigation and Adaptation. This research was also supported by the University of Ljubljana Grant SN-ZRD/22-27/0510. The authors are grateful to Peter Bechtold, Lina Boljka, Massimo Bonavita, Hans Hersbach, Michael Mayer, Dinand Scheppers, and Benoit Vanniere for fruitful discussions on the topic. The authors would like to thank colleagues in the Atmospheric Sciences group at the University of Ljubljana for carefully reading the manuscript.
Open Research
Radiosonde measurements data from the Integrated Global Radiosonde Archive Version 2.2 used for the analysis in this article are available at
https://www.ncei.noaa.gov/data/integrated-global-radiosonde-archive/archive/. The ERA5 datasets are freely available from the open-access Climate Data Store (CDS) Catalogue, operated by the European Centre for Medium-Range Weather Forecasts (ECMWF) within the Copernicus Programme (https://cds.climate.copernicus.eu/#!/home). The JRA-55 reanalysis datasets are freely available from the National Centre for Atmospheric Research (NCAR) Data Archive web page (https://rda.ucar.edu/datasets/). The NOAA 20CRv3 reanalysis datasets are freely available on the National Oceanic and Atmospheric Administration (NOAA) Physical Science Laboratory web page
(https://psl.noaa.gov/data/gridded/data.20thC_ReanV3.html). The MERRA-2 reanalysis datasets are freely available at the https://disc.gsfc.nasa.gov/ web page, managed by the NASA Goddard Earth Sciences (GES) Data and Information Services Center (DISC).
Code availability. The introduced analysis was performed using Python. The codes used in this study will be open-access and available through a designated GitHub profile or an open Zenodo repository.